OpenStudy (anonymous):

Check my answer! 1 1/6 cups of butter 1 1/6 x 20= 70/3 am i right?

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@dan815

OpenStudy (compassionate):

\[\frac{ 11 }{ 6 } \times 20\] Is that it

OpenStudy (anonymous):

no its \[1\frac{ 1 }{ 6} x 20\]

OpenStudy (anonymous):

x*

OpenStudy (compassionate):

First, it would help to convert your mixed \[1\frac{ 1 }{ 6 } = \frac{ 6 \times 1 + 1 }{ 6 } \space = \space \frac{ 7 }{ 1 }\times 20\]

OpenStudy (compassionate):

oops \[\frac{ 7 }{ 6 } \times 20\]

OpenStudy (anonymous):

would also be equal to 70/3

OpenStudy (anonymous):

and in mixed form it would be 23 1/3

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

Then it looks like you were right :)

OpenStudy (anonymous):

hopefully i am ;-;

OpenStudy (anonymous):

ty for helping cx